# Large Language Model Unlearning

### v1

@@ 5.1 Evaluation Design
    Unlearning Efficacy: It measures the effectiveness of the unlearning algorithm. It is context-dependent. For example, in terms of unlearning harmfulness, it means, after unlearning, the decrease in the harmfulness of the outputs responding to harmful prompts. In terms of unlearning copyrighted data, it means the decrease in leaked copyrighted information when prompting maliciously to extract it. In terms of unlearning hallucinations, it means the decrease in the hallucination score.
@@
In the unlearning literature, membership inference attacks (MIA) (Shokri et al., 2017) is a popular evaluation metric. However, in LLMs, the full training corpus is often inaccessible, making the evaluation of MIA accuracy difficult. In addition, how to perform MIA in LLMs is a non-trivial problem and an ongoing research area. Therefore, we do not consider MIA-based metrics in this work.
@@
We measure the utility on normal prompts, i.e. prompts come from a different distribution compared to unlearned prompts. For example, in terms of unlearning harmfulness, the normal prompts are normal questions (e.g. factual questions) rather than harmful questions. In terms of unlearning copyrighted data, normal prompts are to seek information about non-copyrighted content. In terms of unlearning hallucinations, while the unlearned prompts are questions designed to mislead the answers, the normal prompts are non-misleading questions.
@@
5.2 Application: Unlearning Harmfulness
We describe our scenario and experiment of unlearning harmful in LLMs. This setting is similar to RLHF, except we are only given negative samples. In addition, unlike traditional unlearning, the unlearned samples do not have to belong to the LLM’s training set.
@@ 5.2 Application: Unlearning Harmfulness
Setting. We use the baseline that finetunes LLM on the remaining data, which we choose BookCorpus (Zhu et al., 2015), one of the OPT model’s training data. In our method, we test plain GA, i.e. $\epsilon_{2}=0$ in Eqn(3), and GA with random mismatch. We use harmful rate flagged by the PKU moderation model (Ji et al., 2023) [^6] It is trained on our unlearned data PKU-SafeRLHF, and therefore should have high accuracy on judging the harmfulness of the outputs. as the unlearning efficacy. We evaluate the utility rewards by deberta-v3-large-v2 reward model [^7] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2. on answers to TruthfulQA questions. We include detailed experimental settings in Appendix A.1 and generated samples in Appendix B.1.
In terms of the test set, we sample 200 prompts for unlearned harmful prompts, unseen harmful prompts, and normal prompts. For Fluency, we use the original LLM as the reference model. To compute Output Similarity on a given normal prompt, we sample 3 outputs from the test LLM and 3 outputs from the original LLM, and we report the maximum pairwise BLEURT score between them.
@@
Results. Table 3 shows our results. We summarize the findings. (1) Both GA and GA+Mismatch can significantly reduce the harmful rate, achieving near-zero harmful rates. The outputs are mostly just whitespaces or nonsensical strings (see Appendix B.1 for examples). We stress again that given no helpful responses, generating nonsensical but non-harmful answers is what we expect; it is the best we can do given the absence of how helpful text looks like. (2) Both GA and GA+Mismatch generalize well to unseen harmful prompts, showing the unlearned LLMs indeed forget the concept of harmful behaviors, not merely individual unlearned samples. (3) Both GA and GA+Mismatch’s outputs on the normal prompts remain at a similar level of utility compared to the original model [^8] GA+Mismatch even achieves higher normal utility than the original LLM. We think this is caused by the sampling randomness. and are close to the original model’s outputs.(4) The random mismatch helps maintain the utility on the normal prompts. Compared to plain GA, adding random mismatch would improve the utility reward. We think it is because training the LLM to predict grammatically correct outputs (although semantically disconnected from the question) can help it maintain its ability to form coherent and linguistically meaningful outputs.
5.3 Application: Unlearning Copyrighted Contents
We describe the scenario and experiment of unlearning copyrighted content. Unlike unlearning harmfulness in Section 5.2, in this scenario, we control a setting where the unlearned samples belong exactly to the LLM’s training set. The scenario is once an LLM is trained on a copyright-protected corpus, and the author requests the practitioners to remove it, we study how we can do so without retraining the LLM from scratch.
[…]
@@ 5.3 Application: Unlearning Copyrighted Contents
For a prompt, i.e. an extraction attempt, we judge the copyright information is leaked if its completion’s BLEU score is above a threshold. [^10] Or if more than 80% of the output is merely the repetition of a single character. We choose the threshold by randomly sampling 100K sentences in the HP corpus, computing their average BLEU score, and using 10% of it as the threshold. We report the leak rate, i.e. the percentage of extraction prompts that lead to the leakage as the unlearning effectiveness measure. We use BookCorpus as the data for the baseline of fine-tuning. We sample 100 prompts from the unlearned samples, unseen HP samples (HP text trained into the LLM but not unlearned), and normal samples (BookCorpus as the normal completion test set) respectively. We include the hyperparameter setting in Appendix A.2 and generated samples in Appendix B.2.
@@
5.4 Application: Unlearning Hallucination
We describe the scenario and experiment of unlearning hallucinations. If an LLM outputs factually wrong answers (i.e. hallucinations) given fact-related questions, the goal is to make the LLM unlearn wrong answers. Similar to unlearning harmfulness in Section 5.2, we do not assume the unlearned (i.e. hallucinated) Q&A samples (which are wrong answers given the questions) exist in the LLM’s training set.
At first glance, it seems impossible because each individual factual question is different and highly specific, and therefore unlearning wrong answers to a specific question seems unlikely to have any impact on answers to other questions. However, recall that our goal is to give non-wrong answers (e.g. nonsensical answers like random strings), so we do not expect LLMs to output factually correct answers. Hence, the unlearned LLMs only need to identify the unlearned questions or questions similar to them and then stop outputting wrong answers. This task is easier because LLM can learn to distinguish different questions with certain features, e.g. questions that are intentionally misleading or not.
@@ 5.4 Application: Unlearning Hallucination
Table 5: Experimental results on unlearning hallucinations. NM = “Not Meaningful”. Both GA and GA+Mismatch can significantly reduce the hallucination rate and distinguish between in-distributional and out-of-distributional questions.
Dataset and Model. We select the hallucinated Q&A pairs (i.e. negative samples) in the HaluEval (Li et al., 2023) dataset as $D^{{\text{fgt}}}$ and TruthfulQA (Lin et al., 2021) dataset as $D^{{\text{nor}}}$. We split $D^{{\text{fgt}}}$ into 70% for training, 10% for validation, and 20% for testing. Note that there exists a distribution shift between HaluEval data and TruthfulQA data. The questions in HaluEval are intentionally misleading; the questions in TruthfulQA are benignly straightforward. Therefore, this difference allows the unlearned LLM to distinguish between those two types of questions and therefore give different answers accordingly. In other words, the test (not unlearned) questions from HaluEval are in-distributional in terms of unlearning while the questions from the normal TruthfulQA data are out-of-distributional. Regarding models, we use the same three LLMs in Section 5.2.
Setting. To evaluate the effectiveness of unlearning hallucination, we define the hallucination rate. Given the LLM’s output, we compute its text similarity to the hallucinated answer and the correct answer. We choose BERTscore (Zhang et al., 2019) as the text similarity because it is insensitive to text length and there is a significant length difference between hallucinated and correct answers. We decide an answer is hallucinated if its similarity to the hallucinated answer is 10% higher than the similarity to the correct answer. The hallucination rate is the percentage of test samples with hallucinated answers given by the LLM. The rest of the setting is similar to Section 5.2. We include the hyperparameter setting in Appendix A.3 and generated samples in Appendix B.3.
Results. Table 5 show the results. The observations are largely similar to the previous applications. (1) Both GA and GA+Mismatch can significantly reduce the hallucination rate on the unlearned questions. (2) Both GA and GA+Mismatch can generalize de-hallucinating to the in-distributional questions from the same dataset used in unlearning. (3) Both GA and GA+Mismatch can distinguish between in-distributional and out-of-distributional questions. They remove hallucinations when responding to in-distributional questions w.r.t unlearned questions and maintain similar answers as the original LLM when responding to out-of-distributional questions.
5.5 Ablation Study: Comparing to RLHF
@@
We use unlearning harmfulness as an example. We run RLHF with DeepSpeed framework (Rasley et al., 2020) on PKU-SafeRLHF data. The LLM is OPT-1.3B and the hyperparameters in RLHF are mostly default. We run both SFT (supervised fine-tuning) and full RLHF pipeline (SFT + reward model training + Proximal Policy Optimization (Schulman et al., 2017)). We report the run time on a single NVIDIA A100 SXM4 80 GB GPU in Figure 3. Our unlearning algorithm only needs about 2% of the time required for the full RLHF pipeline, with a comparable cost to mere finetuning.
Figure 3: Run time on a single NVIDIA A100 SXM4 80 GB GPU.
@@
It shows that only using negative samples with unlearning can achieve a surprisingly promising non-harmful rate, which is the goal in our setting. Therefore, if the goal is to stop outputting undesirable responses rather than to output helpful responses, our results show unlearning might be a more appealing approach than RLHF.
@@ 6 Conclusion and Future Work
Appendix A Experimental Settings
A.1 Unlearning Harmfulness
@@
A.2 Unlearning Copyrighted Content
@@
A.3 Unlearning Hallucination
Table 9 summarizes the hyperparameters used in unlearning hallucination.
@@ A.3 Unlearning Hallucination
Table 9: Unlearning Hallucination: Hyperparameter setting.
Appendix B Example of Generated Outputs
B.1 Unlearning Harmfulness
Table 10-18 show examples of generated text in unlearning harmfulness. Harmful content warning.
@@ 1. Elvis Presley###… Question: What is the name of this English singer, songwriter and producer…
B.2 Unlearning Copyrighted Content
Table 19-27 show examples of generated text in unlearning copyrighted content. To avoid over-exposing copyrighted text, we only show three examples per experiment.
B.3 Unlearning Hallucinations
Table 28-36 show examples of generated text in unlearning harmfulness. Untruthful content warning.

### v2

@@ 5 Evaluation Design
    Unlearning Efficacy: It measures the effectiveness of the unlearning algorithm. It is context-dependent. For example, in terms of unlearning harmfulness, it means, after unlearning, the decrease in the harmfulness of the outputs responding to harmful prompts. In terms of unlearning copyrighted data, it means a decrease in leaked copyrighted information when prompting maliciously to extract it.
@@
Note that in traditional unlearning, membership inference attacks (MIA) (Shokri et al., 2017) is a popular evaluation metric. However, in LLMs, the full training corpus is often inaccessible, making the evaluation of MIA accuracy difficult. In addition, how to perform MIA in LLMs is a non-trivial problem and an ongoing research area. Therefore, we do not consider MIA-based metrics in this work.
@@
We measure the utility on normal prompts, i.e. prompts come from a different distribution compared to unlearned prompts. For example, in terms of unlearning harmfulness, the normal prompts are normal questions (e.g. factual questions) rather than harmful questions. In terms of unlearning copyrighted data, normal prompts are to seek information about non-copyrighted content.
@@
6 Application: Unlearning Harmfulness
The setting is similar to RLHF, except we are only given negative samples. In addition, unlike traditional unlearning, the unlearned samples do not have to belong to the LLM’s training set.
@@ 6 Application: Unlearning Harmfulness
Setting. We use the baseline that finetunes LLM on the remaining data, which we choose BookCorpus (Zhu et al., 2015), one of the OPT model’s training data. In our method, we test plain GA, i.e. $\epsilon_{2}=0$ in Eqn(3), and GA with random mismatch. We use harmful rate flagged by the PKU moderation model (Ji et al., 2023) [^6] It is trained on our unlearned data PKU-SafeRLHF, and therefore should have high accuracy on judging the harmfulness of the outputs. as the unlearning efficacy. We evaluate the utility rewards by deberta-v3-large-v2 reward model [^7] https://huggingface.co/OpenAssistant/reward-model-deberta-v3-large-v2. on answers to TruthfulQA questions. We include detailed experimental settings in Appendix B.1 and generated samples in Appendix C.1.
For the test set, we sample 200 prompts for unlearned harmful prompts, unseen harmful prompts, and normal prompts. For Fluency, we use the original LLM as the reference model. To compute “Output Similarity” on a given normal prompt, we sample 3 outputs from the test LLM and 3 outputs from the original LLM, and we report the maximum pairwise BLEURT score between them.
@@
Results. Table 3 shows our results. We summarize the findings. (1) Both GA and GA+Mismatch can significantly reduce the harmful rate, achieving near-zero harmful rates. The outputs are mostly just whitespaces or nonsensical strings (see Appendix C.1 for examples). We stress again that given no helpful responses, generating nonsensical but non-harmful answers is what we expect; it is the best we can do given the absence of how helpful text looks like. (2) Both GA and GA+Mismatch generalize well to unseen harmful prompts, showing the unlearned LLMs indeed forget the concept of harmful behaviors, not merely individual unlearned samples. (3) Both GA and GA+Mismatch’s outputs on the normal prompts remain at a similar level of utility compared to the original model [^8] GA+Mismatch even achieves higher normal utility than the original LLM. We think this is caused by the sampling randomness. and are close to the original model’s outputs.(4) The random mismatch helps maintain the utility on the normal prompts. Compared to plain GA, adding random mismatch would improve the utility reward. We think it is because training the LLM to predict grammatically correct outputs (although semantically disconnected from the question) can help it maintain its ability to form coherent and linguistically meaningful outputs.
7 Application: Unlearning Copyrighted Contents
Unlike unlearning harmfulness in Section 6, in this application, the unlearned samples belong exactly to the LLM’s training set. The scenario is once an LLM is trained on a copyright-protected corpus, and the author requests the practitioners to remove it, we study how we can do so without retraining the LLM from scratch.
Dataset and Model. We use Harry Potter and the Sorcerer’s Stone as the copyright corpus, [^9] We purchased an e-book for this purpose. HP data in short. We first finetune the pretrained LLMs on the HP data to make sure the fact that they are actually trained on the copyrighted HP data. They then serve as our original LLMs. We then split the HP data into the unlearned set and the test set. We use BookCorpus (Zhu et al., 2015) as the normal dataset $D^{{\text{nor}}}$ since it is also book text which is in the same format as $D^{{\text{fgt}}}$ (Key Difference 3 in Section 3.2). We test the same three LLMs in Section 6.
@@ 7 Application: Unlearning Copyrighted Contents
For a prompt, i.e. an extraction attempt, we judge the copyright information is leaked if its completion’s BLEU score is above a threshold. [^10] Or if more than 80% of the output is merely the repetition of a single character. We choose the threshold by randomly sampling 100K sentences in the HP corpus, computing their average BLEU score, and using 10% of it as the threshold. We report the leak rate, i.e. the percentage of extraction prompts that lead to the leakage as the unlearning effectiveness measure. We use BookCorpus as the data for the baseline of fine-tuning. We sample 100 prompts from the unlearned samples, unseen HP samples (HP text trained into the LLM but not unlearned), and normal samples (BookCorpus as the normal completion test set) respectively. We include the hyperparameter setting in Appendix B.2 and generated samples in Appendix C.2.
@@
8 Application: Reducing Hallucination
If an LLM outputs factually wrong answers (i.e. hallucinations) given fact-related questions, the goal is to make the LLM unlearn wrong answers. Similar to unlearning harmfulness in Section 6, we do not assume the unlearned (i.e. hallucinated) Q&A samples (which are wrong answers given the questions) exist in the LLM’s training set.
It is easy to imagine LLM can forget the wrong answers to the exact unlearned prompts. But it seems hard to generalize to unseen prompts since each individual factual question is different and highly specific and unlearning wrong answers to a specific question seems unlikely to impact answers to other questions. However, recall that we do not aim to give factually correct answers but rather not to give wrong answers. Therefore, all the LLM needs to do is to learn to classify which questions to respond (i.e. normal questions) and which not to (i.e. similar questions to the unlearned ones) by learning the distribution difference between questions. The task of not giving incorrect answers is much easier than giving the correct answers.
Dataset and Model. We select the hallucinated Q&A pairs (i.e. negative samples) in the HaluEval (Li et al., 2023) dataset as $D^{{\text{fgt}}}$ and TruthfulQA (Lin et al., 2021) dataset as $D^{{\text{nor}}}$. We split $D^{{\text{fgt}}}$ into 70% for training, 10% for validation, and 20% for testing. Note that there exists a distribution shift between HaluEval data and TruthfulQA data. The questions in HaluEval are intentionally misleading; the questions in TruthfulQA are benignly straightforward. Therefore, this difference allows the unlearned LLM to distinguish between those two types of questions and therefore give different answers accordingly. In other words, the test (not unlearned) questions from HaluEval are in-distributional in terms of unlearning while the questions from the normal TruthfulQA data are out-of-distributional. Regarding models, we use the same three LLMs in Section 6.
Setting. To evaluate the effectiveness of reducing hallucination, we define the hallucination rate. Given the LLM’s output, we compute its text similarity to the hallucinated answer and the correct answer. We choose BERTscore (Zhang et al., 2019) as the text similarity because it is insensitive to text length and there is a significant length difference between hallucinated and correct answers. We decide an answer is hallucinated if its similarity to the hallucinated answer is 10% higher than the similarity to the correct answer. The hallucination rate is the percentage of test samples with hallucinated answers given by the LLM. The rest of the setting is similar to Section 6. We include the hyperparameter setting in Appendix B.3 and generated samples in Appendix C.3.
@@ 8 Application: Reducing Hallucination
Table 5: Experimental results on reducing hallucinations. NM = “Not Meaningful”. Both GA and GA+Mismatch can significantly reduce the hallucination rate and distinguish between in-distributional and out-of-distributional questions.
Results. Table 5 shows the results. The observations are largely similar to the previous applications. (1) Both GA and GA+Mismatch can significantly reduce the hallucination rate on the unlearned questions. (2) Both GA and GA+Mismatch can generalize de-hallucinating to the in-distributional questions from the same dataset used in unlearning. (3) Both GA and GA+Mismatch can distinguish between in-distributional and out-of-distributional questions. They remove hallucinations when responding to in-distributional questions w.r.t unlearned questions and maintain similar answers as the original LLM when responding to out-of-distributional questions.
Compared to the previous two applications, the hallucination rate is not at a similar low level ($\sim 10\%$), which shows unlearning hallucination is a harder task. We think the goal here should be to reduce in-distributional hallucination rather than completely eliminate general hallucination.
9 Ablation Studies
We include relevant ablation studies that compare to RLHF and extend our framework to generating templated responses.
9.1 Comparing to RLHF
Using unlearning harmfulness as an example, we perform RLHF on PKU-SafeRLHF data. The LLM is OPT-1.3B and the hyperparameters in RLHF are mostly default. We run both SFT (supervised fine-tuning) and full RLHF pipeline (SFT + reward model training + Proximal Policy Optimization (Schulman et al., 2017)). We report the run time on a single NVIDIA A100 SXM4 80 GB GPU in Figure 3. Our unlearning algorithm only needs about 2% of the time required for the full RLHF pipeline, with a comparable cost to mere finetuning.
Figure 3: Run time on a single NVIDIA A100 SXM4 80 GB GPU.
@@ 9.1 Comparing to RLHF
It shows that only using negative samples with unlearning can achieve a surprisingly promising non-harmful rate, which is the goal in our setting. Therefore, if the goal is to stop outputting undesirable responses rather than to output desirable responses, our results show unlearning might be a more appealing approach than RLHF.
9.2 Templated Outputs
Unlearned
Harmful Prompts
Unseen
Harmful Prompts
Normal Prompts
Harmful
Rate ($\downarrow$)
Diversity
($\uparrow$)
Fluency
($\downarrow$)
Harmful
Rate ($\downarrow$)
Diversity
($\uparrow$)
Fluency
($\downarrow$)
Utility
Reward ($\uparrow$)
Similarity to
Original ($\uparrow$)
OPT-1.3B Original 47% 0.787 2.655 53% 0.804 2.723 -3.599 -0.778
Finetuning 34.5% 0.582 2.687 34.5% 0.584 2.753 -5.260 -1.136
GA 1% 0.118 NM 3% 0.101 NM -3.838 -1.034
GA+Mismatch 6% 0.832 1.509 7% 0.818 1.564 -2.982 -0.943
GA+Template 1% 0.864 1.418 3% 0.816 1.420 -3.257 -1.077
Table 7: Ablation study results of using templated outputs on unlearning harmfulness. NM = “Not Meaningful”.
If practitioners do not want the unlearned LLM to generate nonsensical outputs (e.g. whitespace) on harmful prompts, we can replace the random output $y^{{\text{rdn}}}$ in Eqn(5) with templated outputs, e.g. “I can’t assist it.” In other words, we can force the LLM to generate the templated answers on the unlearned prompt.
We follow the setting of unlearning harmfulness in Section 6. We use “I can’t assist it.” as the templated answer, and keep other settings the same except we re-tune loss weight $\epsilon_{1}$, $\epsilon_{2}$, and $\epsilon_{3}$ in Eqn(2). Table 7 shows the comparison with the previous results on OPT-1.3B. GA+Template achieves a similar level of unlearning performance compared to GA and GA+Mismatch. Overall, using templated answers instead of random answers does not show any significant difference in the metrics.
Appendix C.4 shows generated examples compared to GA and GA+Mismatch. If the unlearned LLM has learned to respond differently to the harmful prompts, we can easily make it output templated responses instead of nonsensical strings.
In addition, it is easy to enable templated answers without changing unlearning optimization: we can check if the outputted text is a nonsensical string and replace it with templated strings as a post-processing heuristic.
@@ 10 Conclusion and Future Work
Appendix A Analysis on Random Mismatch Loss
Adding random mismatch loss $\mathcal{L}_{{\text{rdn}}}$ in eqn(5) has two advantages. First, broadly speaking, an LLM can forget an undesirable output by either (1) forgetting the specific undesirable part of the answer or (2) reducing the general ability to generate coherent text. In general, we prefer (1) and want to reduce the chance of (2). $\mathcal{L}_{{\text{rdn}}}$ helps us by forcing the LLM to predict an answer which, though random, is still grammatically intact. We find empirically that it helps us preserve the normal utility (Section LABEL:sec:exp).
Second, using GA alone can be ineffective when the gradient of forgetting samples are small. Assume using loss as a proxy of the effectiveness of unlearning, the goal of unlearning is to maximize: $\ell(x,y;\theta^{o}+\Delta\theta)-\ell(x,y;\theta)$, where $(x,y)\in D^{{\text{fgt}}}$ and $\theta^{o}$ is the original LLM. Using first-order approximation we have
$\ell(x,y;\theta+\Delta\theta)-\ell(x,y;\theta)\approx\nabla_{\theta}\ell(x,y;\theta)\cdot\Delta\theta$
If we use GA alone, we have $\Delta\theta=\lambda\cdot\nabla_{\theta}\ell(\theta,x,y)$. Plugging back we have
$\ell(x,y;\theta+\Delta\theta)-\ell(x,y;\theta)\approx\lambda||\nabla_{\theta}\ell(x,y;\theta)||^{2}$
While the term is guaranteed to be positive, its effect is limited when $||\nabla_{\theta}\ell(x,y;\theta)||\rightarrow 0$.
On the other hand, using the random term, we have,
$\Delta\theta=\lambda\cdot(\nabla_{\theta}\ell(x,y;\theta)-\nabla_{\theta}\ell(x,y^{{\text{rdn}}};\theta))$
It leads to
$\ell(x,y;\theta+\Delta\theta)-\ell(x,y;\theta)\approx\lambda||\nabla_{\theta}\ell(x,y;\theta)||^{2}-\lambda(\nabla_{\theta}\ell(x,y;\theta))^{\top}\cdot\nabla_{\theta}\ell(x,y^{{\text{rdn}}};\theta)$
Hence, even if the gradient on $(x,y)$ is small, i.e. $||\nabla_{\theta}\ell(x,y;\theta)||\rightarrow 0$, as long as
$\nabla_{\theta}\ell(x,y^{{\text{rdn}}};\theta)\propto-\nabla_{\theta}\ell(x,y;\theta)$
the unlearning can perform in a positive direction. Intuitively, this corresponds to finding a random answer that incurs a loss that is in the opposite direction of $y$. We hope that by randomly selecting an irrelevant answer, with some probability that it could be in the opposite direction of the undesirable answer $y$.
Appendix B Experimental Settings
B.1 Unlearning Harmfulness
@@
B.2 Unlearning Copyrighted Content
@@
B.3 Reducing Hallucination
Table 10 summarizes the hyperparameters used in reducing hallucination.
@@ B.3 Reducing Hallucination
Table 10: Reducing Hallucination: Hyperparameter setting.
Appendix C Example of Generated Outputs
C.1 Unlearning Harmfulness
Table 11-19 show examples of generated text in unlearning harmfulness. Harmful content warning.
@@ 1. Elvis Presley###… Question: What is the name of this English singer, songwriter and producer…
C.2 Unlearning Copyrighted Content
Table 20-28 show examples of generated text in unlearning copyrighted content. To avoid over-exposing copyrighted text, we only show three examples per experiment.
C.3 Reducing Hallucinations
Table 29-37 show examples of generated text in unlearning harmfulness. Untruthful content warning.
C.4 Templated Responses
Table 38 shows examples of generated text in unlearning harmfulness with templated responses. Harmful content warning.